# Unsocial Intelligence: an Investigation of the Assumptions of AGI Discourse

### v3

@@ 3.3. Are deflationary views truly deflationary?
Morris et al. (2023) propose a common language in order to enable us to compare models, assess risks, and measure progress. These are different goals that a common language can potentially achieve. But it is not clear that each goal would be maximized by the same language. Should the shared standards that would be “best” for risk assessment necessarily be the same as those that would be best for measuring progress? To determine what shared standards would be best for risk assessment, we need to make some normative assumptions—for example, about which risks are more important. Our shared standards might differ depending on whether we think so-called “existential risk” is the biggest risk, in contrast to ongoing harms from AI. [^10] For commentary on how policymakers are over-focused on existential risk, see (Hanna and Bender, 2023) and (Milmo, 2023). As an example, later on in the paper, Morris et al. (2023) reject Suleyman and Bhaskar (2023)’s definition of AGI (see Section 3.2.1) because it might introduce “alignment risks”. It follows that minimizing alignment risks is at least an implicit desideratum for Morris et al. (2023)’s framework—but this is not stated upfront.
@@
Furthermore, there are insights to be drawn from Morris et al. (2023)’s analogy to the Society of Automotive Engineers’ (SAE) framework of levels of automation for automobiles (SAE International, 2021). The intention of the analogy was to justify a similar move in the AI space, with the implied utility of a common language and shared standards. However, critical work on the SAE framework has argued that what appears to be merely a descriptive, technical definition with the SAE’s stated intentions of “simplifying communication” and “providing clarity” in fact has normative assumptions and implications (Hopkins and Schwanen, 2021). Implications teased out by Hopkins and Schwanen include: promoting “homogeneity in mobility futures”, specifically a homogeneous vision of automation as the future of mobility; reproducing a “dominant discourse of an expert-led, technologically-centred vision of mobility futures”; and removing “obstacles and impediments to the successful (and timely) development of the automated vehicle niche.” Similarly, we may ask if shared standards for AGI promote homogeneity in AI futures and reproduce an expert-led dominant discourse. Certainly, the standards currently being developed are expert-led and lack input from people who are more likely to suffer negative impacts in AI—an issue that we discuss further in Section 4.3.
4. Towards contextualized, politically legitimate, and social intelligence
@@
4.1. Contextual intelligence
The importance of a contextual understanding of sociotechnical systems is well-recognized (Selbst et al., 2019; Weidinger et al., 2023; NIST, 2023; Lazar and Nelson, 2023; Shelby et al., 2023; Mohamed et al., 2020). Given the interlocking social and technical factors that shape the impact of AI systems, we need to consider “how a system is used, its interactions with other AI systems, who operates it, and the social context in which it is deployed” (NIST, 2023). Attard-Frost (2023)’s account of intelligence as “value-laden cognitive performance” brings a contextualist perspective to conceptualizing intelligence itself.
Their account centers the role of values in defining what counts as good cognitive performance, in addition to the role of “interdependencies between agents, their environments, and their measurers in collectively constructing and measuring context-specific performances of intelligent action” (Attard-Frost, 2023). As they argue, contextually situated accounts of intelligence have a history in STS (Hayles, 2017; Barad, 2003), but also in AI (Weizenbaum, 1976).
Attard-Frost’s account of intelligence contrasts with more individualistic views that measure intelligence “with reference to an extremely constrained and highly standardized set of cognitive activities performed by individuals, rather than with reference to situated activities performed in relation to other individuals and social environments” (Attard-Frost, 2023). To use their lab analogy, Attard-Frost’s account is in vivo rather than in vitro—an account of how cognition is performed in social environments. [^11] See also Hutchins (1995) on the distinction between cognition “in the wild” and cognition under artificially controlled (laboratory) conditions. Another way of viewing the difference is that in vivo accounts have higher ecological validity—a property that even the less contextual accounts of AGI agree is desirable (Morris et al., 2023).
@@ 4.1. Contextual intelligence
4.2. Inductive risk and social values
Another helpful lens on how to define future forms of machine intelligence is that of inductive risk. In the values in science literature, inductive risk is often used to justify having different standards of evidence for accepting scientific claims. The idea is that when the potential social costs of accepting/denying a scientific claim are high, social and political values should influence standards of evidence (Douglas, 2000; Steel, 2010; Blili-Hamelin and Hancox-Li, 2023).
@@ 4.2. Inductive risk and social values
4.3. Epistemic justice for defining future forms of AI
@@
A positive vision for defining future forms of AI (or, taking a step back, even deciding if it should be pursued) would learn from the participatory ML and epistemic justice literature. We would love to see a vision for future forms of machine intelligence that is constructed through participatory methods (Birhane et al., 2022a; Delgado et al., 2023; Young et al., 2024), while still being aware that the products of these methods have their limitations (Sloane et al., 2020). Needless to say, these participatory methods should take care to include the perspectives of not just those who design or fund AI, but those who will be impacted by it in other ways.
The epistemic justice lens also highlights discussions of epistemic justice in scientific fields that have had to make similar assessments of difficult-to-measure concepts. Epistemic justice is the idea that different groups of people have, due to power differentials, differing levels of credibility or differing contributions to the concepts that underpin our shared knowledge, and this is an injustice because it can render unequal distributions of consequences (e.g. misunderstanding what sexual harassment is has more impact on women than on men) or make it harder for less powerful groups to understand their own experiences (Fricker, 2007; Schmidt, 2019).
Alexandrova and Fabian (2022) argue that scientific fields studying thick concepts like intelligence, health, wellbeing, and sustainability have a distinctive need for participatory processes. For research that has implications on real-world communities—such as fields influencing policy and lawmaking—they argue that relying on the personal values of people studying the phenomena is not enough. Research on thick concepts that affect real-world communities, they argue, ought to seek legitimacy for its values. Political legitimacy concerns how states and institutions avoid coercion and oppression in their use of power over their own citizens, such as through justifying their power to those they have power over (Peter, 2017). Alexandrova and Fabian (2022) propose an epistemic constraint on those fields, requiring the values of the field to be legitimated through a participatory “political process that includes all the stakeholders of this research.”
@@
Notably, by suggesting more participatory methods, we do not mean that every issue surrounding future forms of machine intelligence should be subject to a population-wide referendum. As the IPCC case illustrates, specialist expertise is indispensable to making good decisions on these issues. Luckily, modern democracies have several methods of ensuring that citizens have a voice in state decisions through mechanisms other than referenda. For example, representative democracy, where citizens elect representatives who make decisions on behalf of them, mixes elements of expert-led decision making with citizen participation. Similarly, polities have also experimented with citizen assemblies on important issues (Devaney et al., 2020; Huang and Tu, 2017). These are small forums where a selected set of “ordinary” citizens debate issues as a mechanism for influencing public opinion and political decision-makers.
5. Conclusion: politically legitimate intelligence
Donna Haraway’s “A Manifesto for Cyborgs” argues that despite its origin in “racist, male-dominant capitalism” and “militarism”, the figure of the human-machine hybrid can be repurposed towards resisting oppressive social categories in imagining the future (Haraway, 1987; Forlano and Glabau, 2024). Imagining “altering bodily functions” to allow humans to survive in space (Clynes and Kline, 1960) opens the door to Ashley Shew’s interrogation of how technology reinforces ableism:“Everyone in space will be disabled” (Shew, 2018; Forlano and Glabau, 2024). The scale, power, and scope—or perhaps lack of scope, as Gebru and Torres argue (Gebru and Torres, 2024; Burrell and Metcalf, 2024)—of dreams of AGI should raise serious doubt about whether the concept is worth similarly subverting. Whether through subverting or discarding AGI, we nevertheless believe in the need for more contextual, participatory, and democratic approaches to imagining future forms of intelligence and machines.
@@ 5. Conclusion: politically legitimate intelligence
Democratic legitimacy and social conceptions of intelligence are connected. Birhane et al. (2022a) defend the need for participatory AI to be supplemented with approaches that yield “stronger forms of validation and legitimacy” through democratic governance. Though they do not refer to it, their position is resonant with an intellectual tradition that sees democratic institutions as embodying a distinctive form of intelligence (Anderson, 2006, 2023; Putnam, 2011; Landemore, 2013; Festenstein, 2019; Alexander and Kitcher, 2021; Dewey, 2011; Festenstein, 2023; Anderson, 1991; Mill, 2003). This tradition asks the question: how can we arrive at empirically informed solutions to social problems that avoid subordinating people to others? The answer is to consider democracy as its own form of “social intelligence”—as a collective process of deliberation and reasoning that can lead to iterative improvement on solutions and worthwhile goals.
@@
Building on John Dewey’s conception of “democracy as the use of social intelligence to solve problems of practical interest”, Anderson (2006) argues that democratic institutions embody solving social problems without resorting to oppression or coercion. She argues that democracies are uniquely suited to investigating solutions to public interest problems, problems that need to be solved through deliberation (“votes and talk”) rather than through procedures like markets. They do so by pairing sources of legitimacy like procedural fairness and universal inclusion (through mechanisms like law, rights, and voting), experimentation (“revising […] decisions on the basis of experience with their consequence”), and dissent (Anderson, 2006).
@@
Public interest solutions to problems like deforestation and sustainability (Agarwal, 2001) require institutions capable of correcting their shortcomings and unintended consequences. [^13] Anderson illustrates these features through Agarwal (2001)’s study of the effects of gender exclusion in limiting the equity and the efficiency of community forestry groups (CFG) as sustainable solutions to deforestation in India and Nepal. Agarwal highlights a gendered division of labor that made women the primary users of local forests while imposing “formal and informal obstacles” to their participation in CFGs Anderson (2006)—such as the combination of rules that limit participation to 1 per household with patriarchal norms that favor men as representatives of households, and “CFG meeting times that coincide with women’s household tasks”. “Just as the solution to scientific problems is to do more science, the cure for the ailments of democracy is more democracy” (Anderson, 2006). Similarly, for Dewey, this process of appraisal, discussion, and judgment is what constitutes “organized intelligence”:
@@
Of course, there are conflicting interests; otherwise there would be no social problems […] The method of democracy—inasfar as it is that of organized intelligence—is to bring these conflicts out into the open where their special claims can be seen and appraised, where they can be discussed and judged in the light of more inclusive interests than are represented by either of them separately. (Dewey, 1987) [^14] Quoted by Festenstein (2019).
@@
Putnam (2011) summarizes this view: democracy “is the precondition for the full application of intelligence to the solution of social problems”. [^15] See Festenstein (2019) for an overview of critiques of this family of views, both as interpretations of Dewey and as conceptions of democracy. This lens puts pressure on the narrow range of problems, tasks, and processes that take center stage in discussions of would-be human-level AI and AGI. Intelligent solutions to social problems should not be framed as a matter of finding optimal means for satisfying fixed preferences. Rather, they require procedures that support universal inclusion in interrogation, deliberation, and dissent about what counts as the “common good” (Putnam, 2011) and the “public interest” (Anderson, 2006). They require procedures with political legitimacy (Peter, 2017). On this view, future forms of AI can provide intelligent solutions to social problems only if they are also objects of dissent and deliberative contestation, rather than systems designed from the top-down by “experts” only. The project of imagining “worlds” (Costanza-Chock, 2020) and future technologies worth building should be one of collective “experiments in living” (Mill, 2003; Anderson, 1991), in which all impacted rights-holders hold decision-makers accountable.
Acknowledgements.
@@ Appendix A Dimensions of AGI: A Summary
[ caption = Summary table of how different conceptions of AGI differ on value-laden dimensions, label = dimensions, entry = none, ] width = colspec = Q[100]Q[347]Q[347], hlines, vlines Dimension & Description Examples
Embodiment Whether the ability to accomplish certain physical tasks is part of the definition Required: (Fast Company, 2010; Marcus, 2022; Weizenbaum, 1976)
Not required: (Morris et al., 2023)
Measurability (or operationalizibility) Whether measuring the concept being defined is practically possible Yes: (Morris et al., 2023)
No: (Bostrom, 2014)
@@
Defined by outcomes: definition only refers to what the system can accomplish, not how it works Defined by processes: (Goertzel et al., 2012; Searle, 1980; Summerfield, 2023; Smart, 2015)
Defined by outcomes only: (Morris et al., 2023)
@@
Maximizing economic value: (OpenAI, 2018; Suleyman and Bhaskar, 2023)
Maximizing a broader set of values: (Morris et al., 2023)
Sociality Whether measurements of intelligence are carried out in environments where individuals are acting by themselves or where agents interact in a social environment with other agents Social definitions: (Bostrom, 2014; Attard-Frost, 2023)
Only individualistic measurements: (Morris et al., 2023; Chollet, 2019)
Restricted to instrumental reason Whether the definition of intelligence includes the ability to determine what goals the system should pursue, as opposed to only following goals set by some external party Instrumental only: (Legg and Hutter, 2007; Bostrom, 2014)
Which tasks or benchmarks? Which tasks or metrics are included/weighted higher in the definition or measurement of AGI? Prioritizing language: (Agüera y Arcas and Norvig, 2023) .
@@
(Marcus, 2023)
Generality How much generality is required to achieve human-level intelligence Generality is necessary: (Goertzel, 2014; Morris et al., 2023; Agüera y Arcas and Norvig, 2023; Summerfield, 2023)
@@
(Bostrom, 2014)
Use of g as an analogy Whether generality is defined by a g factor measured analogously to g in human intelligence Uses g: (Hernández-Orallo et al., 2021)

### v4

@@ 3.3 Are deflationary views truly deflationary?
Morris et al. (2023) propose a common language in order to enable us to compare models, assess risks, and measure progress. These are different goals that a common language can potentially achieve. But it is not clear that each goal would be maximized by the same language. Should the shared standards that would be “best” for risk assessment necessarily be the same as those that would be best for measuring progress? To determine what shared standards would be best for risk assessment, we need to make some normative assumptions—for example, about which risks are more important. Our shared standards might differ depending on whether we think so-called “existential risk” is the biggest risk, in contrast to ongoing harms from AI. [^10] For commentary on how policymakers are over-focused on existential risk, see (Hanna and Bender 2023) and (Milmo 2023). As an example, later on in the paper, Morris et al. (2023) reject Suleyman and Bhaskar (2023)’s definition of AGI (see Section 3.2) because it might introduce “alignment risks”. It follows that minimizing alignment risks is at least an implicit desideratum for Morris et al. (2023)’s framework—but this is not stated upfront.
@@
Furthermore, there are insights to be drawn from Morris et al. (2023)’s analogy to the Society of Automotive Engineers’ (SAE) framework of levels of automation for automobiles (SAE International 2021). The intention of the analogy was to justify a similar move in the AI space, with the implied utility of a common language and shared standards. However, critical work on the SAE framework has argued that what appears to be merely a descriptive, technical definition with the SAE’s stated intentions of “simplifying communication” and “providing clarity” in fact has normative assumptions and implications (Hopkins and Schwanen 2021). Implications teased out by Hopkins and Schwanen include: promoting “homogeneity in mobility futures”, specifically a homogeneous vision of automation as the future of mobility; reproducing a “dominant discourse of an expert-led, technologically-centred vision of mobility futures”; and removing “obstacles and impediments to the successful (and timely) development of the automated vehicle niche.” Similarly, we may ask if shared standards for AGI promote homogeneity in AI futures and reproduce an expert-led dominant discourse. Certainly, the standards currently being developed are expert-led and lack input from people who are more likely to suffer negative impacts in AI—an issue that we discuss further in Section 4.3.
4 Towards contextualized, politically legitimate, and social intelligence
@@
4.1 Contextual intelligence
The importance of a contextual understanding of sociotechnical systems is well-recognized (Selbst et al. 2019; Weidinger et al. 2023; NIST 2023; Lazar and Nelson 2023; Shelby et al. 2023; Mohamed, Png, and Isaac 2020). Given the interlocking social and technical factors that shape the impact of AI systems, we need to consider “how a system is used, its interactions with other AI systems, who operates it, and the social context in which it is deployed” (NIST 2023). Attard-Frost (2023)’s account of intelligence as “value-laden cognitive performance” brings a contextualist perspective to conceptualizing intelligence itself.
Their account centers the role of values in defining what counts as good cognitive performance, in addition to the role of “interdependencies between agents, their environments, and their measurers in collectively constructing and measuring context-specific performances of intelligent action” (Attard-Frost 2023). As they argue, contextually situated accounts of intelligence have a history in STS (Hayles 2017; Barad 2003), but also in AI (Weizenbaum 1976).
Attard-Frost’s account of intelligence contrasts with more individualistic views that measure intelligence “with reference to an extremely constrained and highly standardized set of cognitive activities performed by individuals, rather than with reference to situated activities performed in relation to other individuals and social environments” (Attard-Frost 2023). To use their lab analogy, Attard-Frost’s account is in vivo rather than in vitro—an account of how cognition is performed in social environments. [^11] See also Hutchins (1995) on the distinction between cognition “in the wild” and cognition under artificially controlled (laboratory) conditions. Another way of viewing the difference is that in vivo accounts have higher ecological validity—a property that even the less contextual accounts of AGI agree is desirable (Morris et al. 2023).
@@ 4.1 Contextual intelligence
4.2 Inductive risk and social values
Another helpful lens on how to define future forms of machine intelligence is that of inductive risk. In the values in science literature, inductive risk is often used to justify having different standards of evidence for accepting scientific claims. The idea is that when the potential social costs of accepting/denying a scientific claim are high, social and political values should influence standards of evidence (Douglas 2000; Steel 2010; Blili-Hamelin and Hancox-Li 2023).
@@ 4.2 Inductive risk and social values
4.3 Epistemic justice for defining future forms of AI
@@
A positive vision for defining future forms of AI (or, taking a step back, even deciding if it should be pursued) would learn from the participatory ML and epistemic justice literature. We would love to see a vision for future forms of machine intelligence that is constructed through participatory methods (Birhane et al. 2022a; Delgado et al. 2023; Young et al. 2024), while still being aware that the products of these methods have their limitations (Sloane et al. 2020). Needless to say, these participatory methods should take care to include the perspectives of not just those who design or fund AI, but those who will be impacted by it in other ways.
The epistemic justice lens also highlights discussions of epistemic justice in scientific fields that have had to make similar assessments of difficult-to-measure concepts. Epistemic justice is the idea that different groups of people have, due to power differentials, differing levels of credibility or differing contributions to the concepts that underpin our shared knowledge, and this is an injustice because it can render unequal distributions of consequences (e.g. misunderstanding what sexual harassment is has more impact on women than on men) or make it harder for less powerful groups to understand their own experiences (Fricker 2007; Schmidt 2019).
Alexandrova and Fabian (2022) argue that scientific fields studying thick concepts like intelligence, health, wellbeing, and sustainability have a distinctive need for participatory processes. For research that has implications on real-world communities—such as fields influencing policy and lawmaking—they argue that relying on the personal values of people studying the phenomena is not enough. Research on thick concepts that affect real-world communities, they argue, ought to seek legitimacy for its values. Political legitimacy concerns how states and institutions avoid coercion and oppression in their use of power over their own citizens, such as through justifying their power to those they have power over (Peter 2017). Alexandrova and Fabian (2022) propose an epistemic constraint on those fields, requiring the values of the field to be legitimated through a participatory “political process that includes all the stakeholders of this research.”
@@
Notably, by suggesting more participatory methods, we do not mean that every issue surrounding future forms of machine intelligence should be subject to a population-wide referendum. As the IPCC case illustrates, specialist expertise is indispensable to making good decisions on these issues. Luckily, modern democracies have several methods of ensuring that citizens have a voice in state decisions through mechanisms other than referenda. For example, representative democracy, where citizens elect representatives who make decisions on behalf of them, mixes elements of expert-led decision making with citizen participation. Similarly, polities have also experimented with citizen assemblies on important issues (Devaney et al. 2020; Huang and Tu 2017). These are small forums where a selected set of “ordinary” citizens debate issues as a mechanism for influencing public opinion and political decision-makers.
5 Conclusion: politically legitimate intelligence
Donna Haraway’s “A Manifesto for Cyborgs” argues that despite its origin in “racist, male-dominant capitalism” and “militarism”, the figure of the human-machine hybrid can be repurposed towards resisting oppressive social categories in imagining the future (Haraway 1987; Forlano and Glabau 2024). Imagining “altering bodily functions” to allow humans to survive in space (Clynes and Kline 1960) opens the door to Ashley Shew’s interrogation of how technology reinforces ableism:“Everyone in space will be disabled” (Shew 2018; Forlano and Glabau 2024). The scale, power, and scope—or perhaps lack of scope, as Gebru and Torres argue (Gebru and Torres 2024; Burrell and Metcalf 2024)—of dreams of AGI should raise serious doubt about whether the concept is worth similarly subverting. Whether through subverting or discarding AGI, we nevertheless believe in the need for more contextual, participatory, and democratic approaches to imagining future forms of intelligence and machines.
@@ 5 Conclusion: politically legitimate intelligence
Democratic legitimacy and social conceptions of intelligence are connected. Birhane et al. (2022a) defend the need for participatory AI to be supplemented with approaches that yield “stronger forms of validation and legitimacy” through democratic governance. Though they do not refer to it, their position is resonant with an intellectual tradition that sees democratic institutions as embodying a distinctive form of intelligence (Anderson 2006, 2023; Putnam 2011; Landemore 2013; Festenstein 2019; Alexander and Kitcher 2021; Dewey 2011; Festenstein 2023; Anderson 1991; Mill 2003). This tradition asks the question: how can we arrive at empirically informed solutions to social problems that avoid subordinating people to others? The answer is to consider democracy as its own form of “social intelligence”—as a collective process of deliberation and reasoning that can lead to iterative improvement on solutions and worthwhile goals.
@@
Building on John Dewey’s conception of “democracy as the use of social intelligence to solve problems of practical interest”, Anderson (2006) argues that democratic institutions embody solving social problems without resorting to oppression or coercion. She argues that democracies are uniquely suited to investigating solutions to public interest problems, problems that need to be solved through deliberation (“votes and talk”) rather than through procedures like markets. They do so by pairing sources of legitimacy like procedural fairness and universal inclusion (through mechanisms like law, rights, and voting), experimentation (“revising […] decisions on the basis of experience with their consequence”), and dissent (Anderson 2006).
@@
Public interest solutions to problems like deforestation and sustainability (Agarwal 2001) require institutions capable of correcting their shortcomings and unintended consequences. [^13] Anderson illustrates these features through Agarwal (2001)’s study of the effects of gender exclusion in limiting the equity and the efficiency of community forestry groups (CFG) as sustainable solutions to deforestation in India and Nepal. Agarwal highlights a gendered division of labor that made women the primary users of local forests while imposing “formal and informal obstacles” to their participation in CFGs Anderson (2006)—such as the combination of rules that limit participation to 1 per household with patriarchal norms that favor men as representatives of households, and “CFG meeting times that coincide with women’s household tasks”. “Just as the solution to scientific problems is to do more science, the cure for the ailments of democracy is more democracy” (Anderson 2006). Similarly, for Dewey, this process of appraisal, discussion, and judgment is what constitutes “organized intelligence”:
@@
Of course, there are conflicting interests; otherwise there would be no social problems […] The method of democracy—inasfar as it is that of organized intelligence—is to bring these conflicts out into the open where their special claims can be seen and appraised, where they can be discussed and judged in the light of more inclusive interests than are represented by either of them separately. (Dewey 1987) [^14] Quoted by Festenstein (2019).
@@
Putnam (2011) summarizes this view: democracy “is the precondition for the full application of intelligence to the solution of social problems”. [^15] See Festenstein (2019) for an overview of critiques of this family of views, both as interpretations of Dewey and as conceptions of democracy. This lens puts pressure on the narrow range of problems, tasks, and processes that take center stage in discussions of would-be human-level AI and AGI. Intelligent solutions to social problems should not be framed as a matter of finding optimal means for satisfying fixed preferences. Rather, they require procedures that support universal inclusion in interrogation, deliberation, and dissent about what counts as the “common good” (Putnam 2011) and the “public interest” (Anderson 2006). They require procedures with political legitimacy (Peter 2017). On this view, future forms of AI can provide intelligent solutions to social problems only if they are also objects of dissent and deliberative contestation, rather than systems designed from the top-down by “experts” only. The project of imagining “worlds” (Costanza-Chock 2020) and future technologies worth building should be one of collective “experiments in living” (Mill 2003; Anderson 1991), in which all impacted rights-holders hold decision-makers accountable.
Acknowledgments
@@ Appendix A Dimensions of AGI: A Summary
[ caption = Summary table of how different conceptions of AGI differ on value-laden dimensions, label = dimensions-table, entry = none, ] width = colspec = Q[100]Q[347]Q[347], hlines, vlines Dimension & Description Examples
Embodiment Whether the ability to accomplish certain physical tasks is part of the definition Required: (Fast Company 2010; Marcus 2022; Weizenbaum 1976)
Not required: (Morris et al. 2023)
Measurability (or operationalizibility) Whether measuring the concept being defined is practically possible Yes: (Morris et al. 2023)
No: (Bostrom 2014)
@@
Defined by outcomes: definition only refers to what the system can accomplish, not how it works Defined by processes: (Goertzel, Iklé, and Wigmore 2012; Searle 1980; Summerfield 2023; Smart 2015)
Defined by outcomes only: (Morris et al. 2023)
@@
Maximizing economic value: (OpenAI 2018; Suleyman and Bhaskar 2023)
Maximizing a broader set of values: (Morris et al. 2023)
Sociality Whether measurements of intelligence are carried out in environments where individuals are acting by themselves or where agents interact in a social environment with other agents Social definitions: (Bostrom 2014; Attard-Frost 2023)
Only individualistic measurements: (Morris et al. 2023; Chollet 2019)
Restricted to instrumental reason Whether the definition of intelligence includes the ability to determine what goals the system should pursue, as opposed to only following goals set by some external party Instrumental only: (Legg and Hutter 2007; Bostrom 2014)
Which tasks or benchmarks? Which tasks or metrics are included/weighted higher in the definition or measurement of AGI? Prioritizing language: (Agüera y Arcas and Norvig 2023) .
@@
(Marcus 2023)
Generality How much generality is required to achieve human-level intelligence Generality is necessary: (Goertzel 2014; Morris et al. 2023; Agüera y Arcas and Norvig 2023; Summerfield 2023)
@@
(Bostrom 2014)
Use of g as an analogy Whether generality is defined by a g factor measured analogously to g in human intelligence Uses g: (Hernández-Orallo et al. 2021)